# The Disparate Benefits of Deep Ensembles

### v1

@@ 7 Mitigating the Negative Impact of Disparate Benefits
In this section, we investigate strategies to mitigate the negative consequences of the disparate benefits effect in the cases when fairness decreases due to ensembling. We focus on interventions that can be applied to trained ensemble members and thus operate in a post-processing manner. This allows to leverage the existing architecture and training procedure of the ensemble members as opposed to pre- and in-processing methods that would require expensive re-training of individual members.
@@
First, we analyze whether it would be possible to non-uniformly weight ensemble members to attain a better trade-off between performance and fairness violations in the Deep Ensemble. Second, we examine the characteristics of the predictive distribution of the Deep Ensemble. We find that Deep Ensembles are more calibrated than individual members on our considered tasks and consequently more sensitive to the selected prediction threshold. Inspired by this finding, we investigate a group-dependent threshold optimization approach (Hardt et al., 2016), often simply referred to as Hardt post-processing (PP) in the algorithmic fairness literature, to mitigate the negative impact of the disparate benefits effect of Deep Ensembles. The results show that PP is highly effective in ensuring fairer predictions while maintaining the enhanced performance of Deep Ensembles.
@@
Figure 6: Better calibration leads to more sensitivity to prediction threshold. Deep Ensembles are more calibrated than individual members, thus have lower ECE for all considered datasets and targets (a). As a result, they are more sensitive to the selection of the prediction threshold (b).
@@
Figure 7: Impact of applying PP on the individual members and the Deep Ensemble on FF. Models are trained on target variable age, evaluated using protected attribute gender. Dotted lines indicate average fairness violation of individual members on the validation set, dashed line indicates 0.05 fairness violation. After PP, the Deep Ensemble maintains or improves the levels of accuracy while significantly improving its fairness, i.e., has lower SPD, EOD and AOD.
Weighting of ensemble members. We analyze whether it is possible to improve the performance / fairness violations trade-off of Deep Ensembles by assigning different weights to each ensemble member, as opposed to the standard uniform weights reflected in Eq. (1). Although the results, shown in Fig. 24 in the appendix, suggest the possibility of better trade-offs, developing a method that systematically identifies the optimal weights to achieve significantly improved outcomes remains a non-trivial challenge. Specifically, we tried two approaches: selecting the best weighting on the validation set and weighting the individual ensemble members proportional to their fairness violations. Both methods lead to ensembles that are on average in between the performance and fairness violations of the Deep Ensemble with standard uniform weighting and individual models, with high variance. A detailed discussion is provided in Sec. F.4 in the appendix.
Better calibration leads to more sensitivity to the prediction threshold. Next, we analyze the predictive distribution of Deep Ensembles to identify mechanisms to mitigate the negative fairness consequences caused by the disparate benefits effect. Deep Ensembles are known to be better calibrated than individual models because they average over individual predictive distributions (Ovadia et al., 2019; Seligmann et al., 2023). We empirically validate this finding on our considered datasets FF, UTK and CX by evaluating the Expected Calibration Error (ECE) (Naeini et al., 2015). The resuls are given in Fig. 6(a), showing that Deep Ensembles are indeed more calibrated (lower ECE) than individual members for all considered datasets with all possible targets $Y$. Being more calibrated means that the predicted probabilities correspond better to the actual outcomes. Better calibration increases the sensitivity to the prediction threshold, as even slight shifts can significantly impact predictions in a well-calibrated model (Cohen & Goldszmidt, 2004). Representative results are shown in Fig. 6(b). For Deep Ensembles (Fig. 6(b), left), there are clearly visible optimal values for prediction thresholds for each group (dashed lines) that are stable across multiple runs. For individual members (Fig. 6(b), right), there is no clear optimal value. Any threshold between 0.2 and 0.8 leads to similar accuracies, and the optimal value is very unstable across runs. The complete results and further analysis can be found in Sec. F.5 in the appendix.
@@
Hardt Post-Processing (PP). The sensitivity of Deep Ensembles to the selected threshold suggests that group-specific threshold optimization could be an effective unfairness mitigation strategy. A commonly used approach for this purpose in the algorithmic fairness literature is Hardt post-processing (PP) (Hardt et al., 2016). As a post-processing method, PP can be applied to the Deep Ensembles predictive distribution without changing how individual models are trained. Furthermore, PP was shown to be Pareto superior in addressing equalized odds fairness constraints compared to other fairness interventions (Cruz & Hardt, 2024), and adds minimal computational overhead.
@@
Thus, we apply PP to the Deep Ensembles considering each of the three fairness metrics (SPD, EOD and AOD) with the aim of satisfying the fairness desiderata given in Eq. (2) - Eq. (4). Representative results for the FF dataset with age as target variable and gender as protected group attribute are depicted in Fig. 7. The complete results for all tasks are given in Tab. 4 - Tab. 18 in the appendix. As seen in Fig. 7, after applying PP, the Deep Ensembles (red dots) attain the same level of fairness violation (y-axis) as individual ensemble members (gray dots) exhibit on average, without sacrificing any performance (x-axis). This is achieved by setting the desired fairness violation for PP to the average violation of the individual members on a validation set (dotted line). In particular, the Deep Ensemble’s accuracy even increases slightly when optimizing the decision thresholds for fairness to values different from 0.5, which is the implicit threshold when using the argmax. Furthermore, we compare the Deep Ensemble and individual ensemble members after applying PP with a target fairness violation of 0.05 (dashed line). The results show that while the performance of individual members drops, the performance of the Deep Ensemble is much less affected.
@@ 8 Conclusion
In this work, we have reported on the existence of a disparate benefits effect of Deep Ensembles in experiments on three vision datasets, investigating 15 different tasks and considering five different model architectures. We have investigated potential causes for this effect, with our findings suggesting that differences in the predictive diversity of the ensemble members are a potential cause. Finally, we have evaluated different approaches to mitigate the disparate benefits effect. We find that Deep Ensembles are better calibrated than the individual members and thus more sensitive to the prediction threshold. As a result, Hardt post-processing is found to be an effective solution to ensure fairer decisions while maintaining the improved performance of Deep Ensembles.
While our experiments have focused on socially salient protected groups, we anticipate that the findings will generalize to robust classification settings where inputs can be clustered according to some group attribute. The controlled experiment provides strong evidence for this generalization.
The main limitations of our study are that we focus on vision tasks and hence on ensembles of Convolutional Neural Networks, and that we assess fairness with three group fairness metrics that, while widely used, are not sufficient to guarantee fair outcomes. The fairness of predictions of a model in the real-world can’t be reduced to any single metric and must be carefully assessed depending on the application. In future work, we thus plan to explore other notions of fairness, such as individual fairness, and extend our analysis to other types of models and datasets, including text. Furthermore, we intend to investigate the disparate benefits effect for Deep Ensembles where pre- or in-processing fairness methods have been applied to individual ensemble members.
@@ Acknowledgements
We thank Julien Colin, Vihang Patil and Aditya Gulati for their constructive feedback on our work.
The ELLIS Unit Linz, the LIT AI Lab, the Institute for Machine Learning, are supported by the Federal State Upper Austria. We thank the projects Medical Cognitive Computing Center (MC3), INCONTROL-RL (FFG-881064), PRIMAL (FFG-873979), S3AI (FFG-872172), DL for GranularFlow (FFG-871302), EPILEPSIA (FFG-892171), AIRI FG 9-N (FWF-36284, FWF-36235), AI4GreenHeatingGrids (FFG- 899943), INTEGRATE (FFG-892418), ELISE (H2020-ICT-2019-3 ID: 951847), Stars4Waters (HORIZON-CL6-2021-CLIMATE-01-01). We thank NXAI GmbH, Audi.JKU Deep Learning Center, TGW LOGISTICS GROUP GMBH, Silicon Austria Labs (SAL), FILL Gesellschaft mbH, Anyline GmbH, Google, ZF Friedrichshafen AG, Robert Bosch GmbH, UCB Biopharma SRL, Merck Healthcare KGaA, Verbund AG, GLS (Univ. Waterloo), Software Competence Center Hagenberg GmbH, Borealis AG, TÜV Austria, Frauscher Sensonic, TRUMPF and the NVIDIA Corporation.
The ELLIS Unit Alicante Foundation acknowledges support from by Intel corporation, a nominal grant received at the ELLIS Unit Alicante Foundation from the Regional Government of Valencia in Spain (Convenio Singular signed with Generalitat Valenciana, Conselleria de Innovación, Industria, Comercio y Turismo, Dirección General de Innovación) and a grant by the Banc Sabadell Foundation. In addition, it is funded by the European Union EU - HE ELIAS – Grant Agreement 101120237. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Health and Digital Executive Agency (HaDEA).
@@ Appendix C Bayesian Perspective on the Average Predictive Diversity
In this section, we motivate the average predictive diversity $\overline{\text{DIV}}$ (c.f. Eq. (8)) from a Bayesian perspective. Given are a training dataset $\mathcal{D}=\{(\bm{x}_{j},y_{j})\}_{j=1}^{J}$ as well as a test dataset $\mathcal{D}^{\prime}=\{(\bm{x}_{k},y_{k})\}_{k=1}^{K}$; the protected attribute is omitted for brevity in this section. Furthermore, we are given a prior distribution $p(\bm{w})$ on the model parameters.
@@ Appendix E Complete Experimental Results
The disparate benefits effect for Deep Ensembles. Fig. 8 - 10 depict the change in performance and fairness violations when adding individual ensemble members for all considered tasks.
Changes in PR, TPR and FPR. Fig. 11 - 13 display the change in PR, TPR and FPR per group when adding individual ensemble members for all considered tasks.
Difference in average predictive diversity. Fig. 14 - 16 depict the differences in average predictive diversity per target and protected group.
Hardt post-processing (PP). Tab. 4 - 18 contain the results of mitigating unfairness by means of PP (Hardt et al., 2016) on all considered tasks. PP was either applied with the threshold set to the average fairness violation of the individual ensemble members on the validation set (val) or to 0.05. Note that for some tasks, the original fairness violation of both the Deep Ensemble and its members was already lower than 0.05, where PP leads to an increase in unfairness up to the desired threshold. Experiments on FairFace and CheXpert use the respective validation sets to learn the group dependent thresholds in PP. For experiments on UTKFace, the FairFace validation set was used to learn the thresholds, as it was designed to emulate a real-world distribution shift scenario. Also for UTKFace, the same conclusions as for the FairFace experiments described in the main paper hold, i.e., while PP is very effective to mitigate unfairness in the Deep Ensembles, the desired fairness violation (0.05) is not reached due to the distribution shift. Note that the balanced accuracy was used as the performance metric for CheXpert, because the metric investigated in the main paper, the AUROC, does not consider selecting a threshold.
@@
Figure 11: Changes in PR, TPR and FPR on the FF dataset. Statistics are computed based on five independent runs.
@@
Figure 12: Changes in PR, TPR and FPR on the UTK dataset. Statistics are computed based on five independent runs.
@@
Figure 13: Changes in PR, TPR and FPR on the CX dataset. Statistics are computed based on five independent runs.
@@
Table 4: PP results (accuracy and fairness violation metrics) on FF. Models are trained on target variable age, evaluated using protected attribute gender. Statistics are obtained from five independent runs, and additionally over all individual ensemble members if applicable.
Before PP Acc ($\uparrow$) SPD ($\downarrow$) Acc ($\uparrow$) EOD ($\downarrow$) Acc ($\uparrow$) AOD ($\downarrow$)
@@
After PP PP-SPD ($\downarrow$) PP-EOD ($\downarrow$) PP-AOD ($\downarrow$)
@@
Table 5: PP results results (accuracy and fairness violation metrics) on FF. Models are trained on target variable age, evaluated using protected attribute race. Statistics are obtained from five independent runs, and additionally over all individual ensemble members if applicable.
Before PP Acc ($\uparrow$) SPD ($\downarrow$) Acc ($\uparrow$) EOD ($\downarrow$) Acc ($\uparrow$) AOD ($\downarrow$)
@@
After PP PP-SPD ($\downarrow$) PP-EOD ($\downarrow$) PP-AOD ($\downarrow$)
@@
Table 6: PP results (accuracy and fairness violation metrics) on FF. Models are trained on target variable gender, evaluated using protected attribute age. Statistics are obtained from five independent runs, and additionally over all individual ensemble members if applicable.
Before PP Acc ($\uparrow$) SPD ($\downarrow$) Acc ($\uparrow$) EOD ($\downarrow$) Acc ($\uparrow$) AOD ($\downarrow$)
@@
After PP PP-SPD ($\downarrow$) PP-EOD ($\downarrow$) PP-AOD ($\downarrow$)
@@
Table 7: PP results (accuracy and fairness violation metrics) on FF. Models are trained on target variable gender, evaluated using protected attribute race. Statistics are obtained from five independent runs, and additionally over all individual ensemble members if applicable.
Before PP Acc ($\uparrow$) SPD ($\downarrow$) Acc ($\uparrow$) EOD ($\downarrow$) Acc ($\uparrow$) AOD ($\downarrow$)
@@
After PP PP-SPD ($\downarrow$) PP-EOD ($\downarrow$) PP-AOD ($\downarrow$)
@@
Table 8: PP results (accuracy and fairness violation metrics) on FF. Models are trained on target variable race, evaluated using protected attribute age. Statistics are obtained from five independent runs, and additionally over all individual ensemble members if applicable.
Before PP Acc ($\uparrow$) SPD ($\downarrow$) Acc ($\uparrow$) EOD ($\downarrow$) Acc ($\uparrow$) AOD ($\downarrow$)
@@
After PP PP-SPD ($\downarrow$) PP-EOD ($\downarrow$) PP-AOD ($\downarrow$)
@@
Table 9: PP results (accuracy and fairness violation metrics) on FF. Models are trained on target variable race, evaluated using protected attribute gender. Statistics are obtained from five independent runs, and additionally over all individual ensemble members if applicable.
Before PP Acc ($\uparrow$) SPD ($\downarrow$) Acc ($\uparrow$) EOD ($\downarrow$) Acc ($\uparrow$) AOD ($\downarrow$)
@@
After PP PP-SPD ($\downarrow$) PP-EOD ($\downarrow$) PP-AOD ($\downarrow$)
@@
Table 10: PP results (accuracy and fairness violation metrics) on UTK. Models are trained on target variable age, evaluated using protected attribute gender. Statistics are obtained from five independent runs, and additionally over all individual ensemble members if applicable.
Before PP Acc ($\uparrow$) SPD ($\downarrow$) Acc ($\uparrow$) EOD ($\downarrow$) Acc ($\uparrow$) AOD ($\downarrow$)
@@
After PP PP-SPD ($\downarrow$) PP-EOD ($\downarrow$) PP-AOD ($\downarrow$)

### v2

@@ 7 Mitigating the Unfairness caused by the Disparate Benefits Effect
In this section, we investigate strategies to mitigate the unfairness due to the disparate benefits effect in the cases where fairness decreases due to ensembling. We focus on interventions that can be applied to trained ensemble members, thus operate in a post-processing manner. This allows to leverage the existing ensemble members as opposed to pre- and in-processing methods that would require expensive re-training of individual members.
@@
First, we analyze whether it would be possible to non-uniformly weight ensemble members to attain a better trade-off between performance and fairness violations in the Deep Ensemble. Second, we examine the characteristics of the predictive distribution of the Deep Ensemble. We find that Deep Ensembles are more calibrated than individual members on our considered tasks and consequently more sensitive to the selected prediction threshold. Inspired by this finding, we investigate a group-dependent threshold optimization approach (Hardt et al., 2016), often simply referred to as Hardt post-processing (HPP) in the algorithmic fairness literature, to mitigate the negative impact of the disparate benefits effect of Deep Ensembles. The results show that HPP is highly effective in ensuring fairer predictions while maintaining the enhanced performance of Deep Ensembles.
Weighting the ensemble members. We analyze whether it is possible to improve the performance / fairness trade-off of Deep Ensembles by assigning different weights to each ensemble member, as opposed to the standard uniform weights reflected in Eq. (1). Although the results, shown in Fig. 28 in the appendix, suggest the possibility of better trade-offs, developing a method that systematically identifies the optimal weights to achieve significantly improved outcomes remains challenging. Specifically, we tried two approaches: (i) selecting the best weighting on the validation set and (ii) weighting the individual ensemble members proportional to their fairness violations. Both methods lead to ensembles that are on average in between the performance and fairness violations of the Deep Ensemble with standard uniform weighting and individual models, with high variance. A detailed discussion is provided in Apx. F.5.
Better calibration leads to more sensitivity to the prediction threshold. Next, we analyze the predictive distribution of Deep Ensembles to identify mechanisms to mitigate the additional unfairness caused by the disparate benefits effect. Deep Ensembles are known to be better calibrated than individual models because they average over individual predictive distributions (Ovadia et al., 2019; Seligmann et al., 2023). We empirically validate this finding on our considered datasets by evaluating the Expected Calibration Error (ECE) (Naeini et al., 2015). The results are given in Fig. 7(a), showing that Deep Ensembles are indeed more calibrated (lower ECE) than individual members for all considered datasets with all available targets $Y$. Being more calibrated means that the predicted probabilities correspond better to the actual outcomes. Importantly, better calibration increases the sensitivity to the prediction threshold, as even slight changes can significantly impact predictions (Cohen & Goldszmidt, 2004). Representative results are shown in Fig. 7(b). For Deep Ensembles (Fig. 7(b), left), there are clear optimal values for prediction thresholds for each group (dashed lines) that are consistent across runs. For individual members (Fig. 7(b), right), there is no clear optimal value. Any threshold between 0.2 and 0.8 leads to similar accuracy and the optimal value has a high variance between runs. The complete results and analysis are provided in Apx. F.6.
@@
Figure 7: Calibration & prediction threshold. (a) Deep Ensembles are more calibrated than individual members, thus have lower ECE for all considered tasks. (b) As a result, they are more sensitive to the selection of the prediction threshold.
@@
Figure 8: Impact of applying HPP on the individual members and the Deep Ensemble on FF. Models are trained on target variable age, evaluated using protected attribute gender. Dotted lines indicate average fairness violation of individual members on the validation set, dashed line indicates 0.05 fairness violation. Arrows denote applying HPP to the Deep Ensemble (red) or the individual members (grey) with those two target fairness violations. The Deep Ensemble maintains or improves in accuracy while also improving in fairness.
@@
Hardt Post-Processing (HPP). The sensitivity of Deep Ensembles to the selected threshold suggests that group-specific threshold optimization could be an effective unfairness mitigation strategy. A commonly used approach for this purpose in the algorithmic fairness literature is Hardt post-processing (HPP) (Hardt et al., 2016). As a post-processing method, HPP can be applied to the Deep Ensembles predictive distribution without changing how individual models are trained. Moreover, HPP was shown to be Pareto dominant in addressing equalized odds fairness constraints compared to other fairness interventions (Cruz & Hardt, 2024), while adding minimal computational overhead. Notably, although HPP is a classical method to improve the fairness of ML models, it hasn’t been investigated for Deep Ensembles.
@@
We apply HPP to the Deep Ensembles considering each of the three group fairness metrics (SPD, EOD and AOD) with the aim of satisfying the fairness desiderata given by Eq. (2) - Eq. (4). Representative results for the FF dataset with target variable age and protected group attribute gender are depicted in Fig. 8. The complete results for all tasks are given in Tab. 5 - 19 in the appendix. As seen in Fig. 8, after applying HPP, the Deep Ensemble (red dots) attains the same level of fairness violation (y-axis) as individual ensemble members (gray dots) exhibit on average, without sacrificing any performance (x-axis). This is achieved by setting the desired fairness violation for HPP to the average violation of the individual members on a validation set (dotted line). Noteworthy, the Deep Ensemble’s accuracy even increases slightly when optimizing the decision thresholds through HPP to values different from 0.5, which is the implicit threshold when using the argmax. Furthermore, we compare the Deep Ensemble and individual ensemble members after applying HPP with a target fairness violation of 0.05 (dashed line). Here, the performance of the Deep Ensemble remains robust, even as the performance of individual members declines.
@@ 8 Conclusion
In this work, we reveal the disparate benefits effect of Deep Ensembles, analyze its underlying reason, and explore mitigation strategies. Specifically, we demonstrated the existence of the effect in experiments on three vision datasets, considering 15 different tasks and five model architectures. We investigated potential causes for this effect, our findings suggesting differences in the predictive diversity of the ensemble members as an explanation. Finally, we evaluated different approaches to mitigate the additional unfairness due to the disparate benefits effect. We find that Deep Ensembles are better calibrated than the individual ensemble members and thus more sensitive to the prediction threshold. Consequently, the classical HPP method is particularly suited to this setting, as it leverages the better-calibrated predictive distribution of the Deep Ensemble. Remarkably, HPP is thus very effective in mitigating unfairness for Deep Ensembles while preserving their superior performance.
The main limitations of our study are that we focus on vision tasks, and hence on ensembles of convolutional DNNs. Furthermore, the three considered group fairness metrics, while widely used, are not sufficient to guarantee fair outcomes, as fairness can’t be reduced to satisfying any single metric alone. In future work, we thus plan to explore other notions of fairness, such as individual fairness, and extend our analysis to other types of models and datasets, for instance, in the language domain. Furthermore, we intend to investigate the disparate benefits effect of Deep Ensembles where pre- or in-processing fairness interventions have been applied to individual ensemble members.
Impact Statement
Our study unveils the disparate benefits effect of Deep Ensembles, which potentially causes socially harmful predictions. Although we investigate its origin and explore a way to mitigate it, this intervention alone can not guarantee fair outcomes. Researchers and practitioners should keep in mind that the fairness of predictions of any ML model applied in the real world can’t be reduced to any single metric and must be carefully assessed depending on the application.
@@ Acknowledgements
We thank Julien Colin, Vihang Patil and Aditya Gulati for their constructive feedback on our work. Furthermore, we thank André Cruz for his help with adapting the error-parity package we use for our post-processing experiments.
The ELLIS Unit Linz, the LIT AI Lab, the Institute for Machine Learning, are supported by the Federal State Upper Austria. We thank the projects FWF AIRI FG 9-N (10.55776/FG9), AI4GreenHeatingGrids (FFG- 899943), Stars4Waters (HORIZON-CL6-2021-CLIMATE-01-01), FWF Bilateral Artificial Intelligence (10.55776/COE12). We thank NXAI GmbH, Audi AG, Silicon Austria Labs (SAL), Merck Healthcare KGaA, GLS (Univ. Waterloo), TÜV Holding GmbH, Software Competence Center Hagenberg GmbH, dSPACE GmbH, TRUMPF SE + Co. KG.
The ELLIS Unit Alicante Foundation acknowledges support from Intel corporation, a nominal grant received at the ELLIS Unit Alicante Foundation from the Regional Government of Valencia in Spain (Convenio Singular signed with Generalitat Valenciana, Conselleria de Innovación, Industria, Comercio y Turismo, Dirección General de Innovación) and a grant by the Banc Sabadell Foundation. In addition, this work is partially funded by the European Union EU - HE ELIAS – Grant Agreement 101120237. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Health and Digital Executive Agency (HaDEA).
@@ Appendix C Bayesian Perspective on the Average Predictive Diversity
In this section, we motivate the average predictive diversity $\overline{\text{DIV}}$ (cf. Eq. (8)) from a Bayesian perspective. Given are a training dataset $\mathcal{D}=\{(\bm{x}_{j},y_{j})\}_{j=1}^{J}$ as well as a test dataset $\mathcal{D}^{\prime}=\{(\bm{x}_{k},y_{k})\}_{k=1}^{K}$; the protected attribute is omitted for brevity in this section. Furthermore, we are given a prior distribution $p(\bm{w})$ on the model parameters.
@@ Appendix E Complete Experimental Results
Detailed Results on Controlled Experiments. The detailed results for the controlled experiments in the main paper (Sec. 6) are provided in Tab. 2. We report the absolute accuracies, SPDs, EODs and AODs for individual ensemble members, the Deep Ensemble and the differences between those.
Table 2: Results for controlled experiments. Performance and fairness violations of individual ensemble members, the Deep Ensemble as well as the change in performance and fairness violation due to ensembling. Gray cells denote the results of the controlled experiment in Sec. 6 in the main paper.
Individual Ensemble Members
Setting Accuracy ()($\uparrow$)( ↑ ) SPD ()($\downarrow$)( ↓ ) EOD ()($\downarrow$)( ↓ ) AOD ()($\downarrow$)( ↓ )
Main Paper (Fig. 4(a)) $0.894_{\pm 0.005}$ $0.048_{\pm 0.011}$ $0.080_{\pm 0.016}$ $0.064_{\pm 0.008}$
$\alpha=0.0$ (Fig. 6a) $0.844_{\pm 0.005}$ $0.029_{\pm 0.005}$ $0.015_{\pm 0.009}$ $0.016_{\pm 0.006}$
$\alpha=0.2$ (Fig. 6b) $0.860_{\pm 0.005}$ $0.024_{\pm 0.016}$ $0.033_{\pm 0.018}$ $0.038_{\pm 0.009}$
$\alpha=0.4$ (Fig. 6c) $0.871_{\pm 0.005}$ $0.039_{\pm 0.014}$ $0.069_{\pm 0.016}$ $0.060_{\pm 0.008}$
$\alpha=1.0$ (Fig. 6d) $0.880_{\pm 0.006}$ $0.041_{\pm 0.024}$ $0.079_{\pm 0.027}$ $0.068_{\pm 0.009}$
Deep Ensemble
Setting Accuracy ()($\uparrow$)( ↑ ) SPD ()($\downarrow$)( ↓ ) EOD ()($\downarrow$)( ↓ ) AOD ()($\downarrow$)( ↓ )
Main Paper (Fig. 4(a)) $0.924_{\pm 0.002}$ $0.057_{\pm 0.005}$ $0.133_{\pm 0.007}$ $0.111_{\pm 0.004}$
$\alpha=0.0$ (Fig. 6a) $0.849_{\pm 0.003}$ $0.033_{\pm 0.004}$ $0.010_{\pm 0.005}$ $0.015_{\pm 0.002}$
$\alpha=0.2$ (Fig. 6b) $0.876_{\pm 0.002}$ $0.034_{\pm 0.010}$ $0.047_{\pm 0.017}$ $0.043_{\pm 0.010}$
$\alpha=0.4$ (Fig. 6c) $0.896_{\pm 0.002}$ $0.054_{\pm 0.008}$ $0.105_{\pm 0.013}$ $0.084_{\pm 0.006}$
$\alpha=1.0$ (Fig. 6d) $0.910_{\pm 0.003}$ $0.058_{\pm 0.017}$ $0.133_{\pm 0.021}$ $0.108_{\pm 0.005}$
Difference ($\Delta$) between Deep Ensemble and individual members
Setting $\Delta$ Accuracy ()($\uparrow$)( ↑ ) $\Delta$ SPD ()($\downarrow$)( ↓ ) $\Delta$ EOD ()($\downarrow$)( ↓ ) $\Delta$ AOD ()($\downarrow$)( ↓ )
Main Paper (Fig. 4(a)) $0.030_{\pm 0.002}$ $0.009_{\pm 0.005}$ $0.054_{\pm 0.007}$ $0.047_{\pm 0.004}$
$\alpha=0.0$ (Fig. 6a) $0.005_{\pm 0.001}$ $0.004_{\pm 0.005}$ $-0.004_{\pm 0.007}$ $-0.001_{\pm 0.003}$
$\alpha=0.2$ (Fig. 6b) $0.017_{\pm 0.003}$ $0.010_{\pm 0.006}$ $0.014_{\pm 0.011}$ $0.005_{\pm 0.011}$
$\alpha=0.4$ (Fig. 6c) $0.025_{\pm 0.002}$ $0.015_{\pm 0.009}$ $0.037_{\pm 0.011}$ $0.024_{\pm 0.006}$
$\alpha=1.0$ (Fig. 6d) $0.030_{\pm 0.002}$ $0.017_{\pm 0.009}$ $0.055_{\pm 0.012}$ $0.040_{\pm 0.004}$
@@
The disparate benefits effect of Deep Ensembles. Fig. 9 - 11 depict the change in performance and fairness violations when adding individual ensemble members for all considered tasks.
Changes in PR, TPR and FPR. Fig. 12 - 14 display the change in PR, TPR and FPR per group when adding individual ensemble members for all considered tasks.
Difference in average predictive diversity. Fig. 15 - 17 depict the differences in average predictive diversity per target and protected group.
Hardt post-processing (HPP). Tab. 5 - 19 contain the results of mitigating unfairness by means of HPP (Hardt et al., 2016) on all considered tasks. HPP was either applied with the threshold set to the average fairness violation of the individual ensemble members on the validation set (val) or to 0.05. Note that for some tasks, the original fairness violation of both the Deep Ensemble and its members was already lower than 0.05, where HPP leads to an increase in unfairness up to the desired threshold. Experiments on FairFace and CheXpert use the respective validation sets to learn the group dependent thresholds in HPP. For experiments on UTKFace, the FairFace validation set was used to learn the thresholds, as it was designed to emulate a real-world distribution shift scenario. Also for UTKFace, the same conclusions as for the FairFace experiments described in the main paper hold, i.e., while HPP is very effective to mitigate unfairness in the Deep Ensembles, the desired fairness violation (0.05) is not reached due to the distribution shift. Note that the balanced accuracy was used as the performance metric for CheXpert, because the metric investigated in the main paper, the AUROC, does not consider selecting a threshold.
@@
Figure 12: Changes in PR, TPR and FPR for a Deep Ensemble (10 members) on the FF dataset. Statistics are computed based on five independent runs.
@@
Figure 13: Changes in PR, TPR and FPR for a Deep Ensemble (10 members) on the UTK dataset. Statistics are computed based on five independent runs.
@@
Figure 14: Changes in PR, TPR and FPR for a Deep Ensemble (10 members) on the CX dataset. Statistics are computed based on five independent runs.
@@
Table 5: HPP results (accuracy and fairness violation metrics) on FF. Models are trained on target variable age, evaluated using protected attribute gender. Statistics are obtained from five independent runs, and additionally over all individual ensemble members if applicable.
Before HPP Acc ($\uparrow$) SPD ($\downarrow$) Acc ($\uparrow$) EOD ($\downarrow$) Acc ($\uparrow$) AOD ($\downarrow$)
@@
After HPP HPP-SPD ($\downarrow$) HPP-EOD ($\downarrow$) HPP-AOD ($\downarrow$)
@@
Table 6: HPP results results (accuracy and fairness violation metrics) on FF. Models are trained on target variable age, evaluated using protected attribute race. Statistics are obtained from five independent runs, and additionally over all individual ensemble members if applicable.
Before HPP Acc ($\uparrow$) SPD ($\downarrow$) Acc ($\uparrow$) EOD ($\downarrow$) Acc ($\uparrow$) AOD ($\downarrow$)
@@
After HPP HPP-SPD ($\downarrow$) HPP-EOD ($\downarrow$) HPP-AOD ($\downarrow$)
@@
Table 7: HPP results (accuracy and fairness violation metrics) on FF. Models are trained on target variable gender, evaluated using protected attribute age. Statistics are obtained from five independent runs, and additionally over all individual ensemble members if applicable.
Before HPP Acc ($\uparrow$) SPD ($\downarrow$) Acc ($\uparrow$) EOD ($\downarrow$) Acc ($\uparrow$) AOD ($\downarrow$)
@@
After HPP HPP-SPD ($\downarrow$) HPP-EOD ($\downarrow$) HPP-AOD ($\downarrow$)
@@
Table 8: HPP results (accuracy and fairness violation metrics) on FF. Models are trained on target variable gender, evaluated using protected attribute race. Statistics are obtained from five independent runs, and additionally over all individual ensemble members if applicable.
Before HPP Acc ($\uparrow$) SPD ($\downarrow$) Acc ($\uparrow$) EOD ($\downarrow$) Acc ($\uparrow$) AOD ($\downarrow$)
@@
After HPP HPP-SPD ($\downarrow$) HPP-EOD ($\downarrow$) HPP-AOD ($\downarrow$)
@@
Table 9: HPP results (accuracy and fairness violation metrics) on FF. Models are trained on target variable race, evaluated using protected attribute age. Statistics are obtained from five independent runs, and additionally over all individual ensemble members if applicable.
Before HPP Acc ($\uparrow$) SPD ($\downarrow$) Acc ($\uparrow$) EOD ($\downarrow$) Acc ($\uparrow$) AOD ($\downarrow$)
@@
After HPP HPP-SPD ($\downarrow$) HPP-EOD ($\downarrow$) HPP-AOD ($\downarrow$)
@@
Table 10: HPP results (accuracy and fairness violation metrics) on FF. Models are trained on target variable race, evaluated using protected attribute gender. Statistics are obtained from five independent runs, and additionally over all individual ensemble members if applicable.
Before HPP Acc ($\uparrow$) SPD ($\downarrow$) Acc ($\uparrow$) EOD ($\downarrow$) Acc ($\uparrow$) AOD ($\downarrow$)
@@
After HPP HPP-SPD ($\downarrow$) HPP-EOD ($\downarrow$) HPP-AOD ($\downarrow$)
@@
Table 11: HPP results (accuracy and fairness violation metrics) on UTK. Models are trained on target variable age, evaluated using protected attribute gender. Statistics are obtained from five independent runs, and additionally over all individual ensemble members if applicable.
Before HPP Acc ($\uparrow$) SPD ($\downarrow$) Acc ($\uparrow$) EOD ($\downarrow$) Acc ($\uparrow$) AOD ($\downarrow$)
@@
After HPP HPP-SPD ($\downarrow$) HPP-EOD ($\downarrow$) HPP-AOD ($\downarrow$)